OpenStudy (anonymous):

Write the following without using logarithms: ln a - ln b + xy = 0 (That's ln of a and ln of b)

OpenStudy (anonymous):

ln a - ln b = ln(a/b) xy = -ln(a/b) = ln(b/a)

OpenStudy (anonymous):

exponentiate both sides

OpenStudy (anonymous):

do I do e^(ln(a/b))=e^(xy) @Algebraic!

hartnn (hartnn):

yes, then use \(\huge e^{ln x}=x\)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^